OpenStudy (anonymous):

solve - 2over3 h = 8over9 will give medal to best answer.

OpenStudy (anonymous):

4over3

OpenStudy (gorv):

2/3 h=8/9 h=8/9 *3/2=4/3

OpenStudy (anonymous):

thanks!